The consumer's utility is given by u(x,y)=16x-3√√x+y. The price of good 2 is normalized to 1. Find the change in consumer's surplus if the price of good 1 changed from 10 to 5.



Answer: 18.35

1
Expert's answer
2021-11-25T19:56:48-0500

"u(x,y)=16x-3\sqrt {\sqrt (x+y)}"

Assume that the consumer initially consumes 10 units of x and 0 units of y at the prices of x being 10 and y being p.

Therefore:

"u(x,y)_{1}=16\times10-3\sqrt {\sqrt10}= 155.59"


When price of x declines to 5 (by half), the consumer can increase his consumption of y by an additional 50 units while reducing their consumption ofx by one units.

Therefore:

"u(x,y)_2=16\times9-3\sqrt {\sqrt69}= 137.24"

"Change in utility surplus =u(x,y)_1-u(x,y)_2"

"Change in utility surplus=155.59-137.24=18.35"